# Exploring Parent’s Needs for Children-Centered AI to Support Preschoolers’ Storytelling and Reading Activities

### v1

@@ 4.1.2. Can AI Replace Parents’ Roles in Storytelling?
4.2. Perceived Value of Using AI-Based Storytelling Technologies
When reflecting on the experience of using AI-based storytelling tools to associate daily parent-child storytelling, our participants talked about some benefits of these technologies they perceived, which will be presented in this section.
To be specific, most of them expressed that they completely agreed that interactive storytelling was a very important and beneficial parent-child activity for preschoolers. Yet, for most of them, this conversation-rich interactive storytelling with preschoolers was full of practical challenges. For example, P9, a mother with a 6-year-old son, said, “I really want to interact with my son through expanding the story content, but I don’t know what content should be expanded. I even can’t understand the questions he asks me.” Some participants also noted that engaging in daily, high-quality interactive storytelling, as they expected, represented significant challenges in terms of time, energy, and physical stamina. This was particularly true for today’s young parents who were often immersed in fast-paced and demanding work.
@@ 4.2. Perceived Value of Using AI-Based Storytelling Technologies
At the same time, many of our participants stated that compared to plush toys and the like, these AI-based tools were relatively “better” toys for children, because they could bring “novelties and a sense of technology” [P8, P13] to children, which parents couldn’t bring. Most existing storytelling tools had some children-tailored interactive designs, such as immersive storytelling processes that blend multiple modalities (“Children could touch, hear, see, and talk with it.” [P6]), and vivid presentations (“The sound of the storyteller is very pleasant and full of emotions and varied intonations, and the animals can move and make sounds” [P1]), which worked well to attract children to participate in the interaction, improving their storytelling experience significantly.
In addition, most of these technologies integrated more or less learning resources, or had some targeted interactive designs, which allowed children to learn something or develop certain abilities while playing. As P8, a mother of 6-year-old son, said: “You always need to buy something for the child, right? Do you buy a bunch of Ultraman or cloth dolls, or these intelligent things? No matter how the practical effect is, it provides the possible opportunities to cultivate the creative thinking abilities of children, or learn something when playing.” [P8]
Despite these perceived benefits, our participants considered that most of the storytelling tools they had bought and used were hard to consistently engage preschoolers and often struggled to achieve the intelligent features claimed by the products. The practical effectiveness and efficiency in supporting quality storytelling were limited. Our participants reported several challenges with these tools that hindered their practical effectiveness. We generally categorized these challenges into two types - interactive and algorithmic, and detailed them in the following two sections.
4.3. Interactive Challenges in Supporting Storytelling Aligned with Cognitive Characteristics
The first category of challenges we identified was interactive challenges, mainly referring to challenges and dilemmas caused by interactive ways and interfaces existing AI-based storytelling tools used, which primarily arose from existing tools’ insufficient consideration of preschool-age children’s cognitive characteristics and their limited abilities to accept and use these technologies.
@@ 4.3.1. Preschoolers are Difficult to Use AI Tools Proactively, Independently, and Consistently
According to our participants, preschoolers often exhibited short attention spans, lacked self-learning and self-management abilities, etc. These characteristics made it hard for them to use various AI-based tools actively, independently, and consistently. Most of the participants thought it was unrealistic to expect preschoolers to use these tools to read a storybook independently. Instead, parents’ engagement was crucial in determining the practical effectiveness of these tools.
To be specific, most existing AI-based storytelling tools require specific usage processes and scenarios. For example, the use of the interactive talking pen (C1 in Table 2) required parents to “take out the pen, turn it on, find the corresponding book, and flip to the specific story [P8]” and then the children could use it to read a story. “Without parents’ operations and guidance, preschoolers are unlikely to proactively use these tools to read the story” [P8]. As for the interactive and question-answering function, our participants expressed that when their children had questions or hoped to interact, asking and interacting with parents was an easier and more straightforward way for them.
Further, preschoolers’ short attention spans made it challenging to consistently and sustainably use any tools with focused attention. Consequently, whether and how parents were involved directly affected the practical effectiveness of tools. Many parents in our study expressed that there was a notable difference in effectiveness between children using these tools independently and when accompanied by parents. P5, a mother of a 6-year-old daughter, considered that, “if parents are not involved, even for the best product, it is difficult to sustain children’s interest and attention, and motivate them to read a story with it for 15-20 minutes steadily”. P8, who had purchased many AI-based storytelling tools for her son, shared her experience with these tools.
Our (parents’) involvement or not has a completely different using effect. I can let these tools tell stories to my son. Tools, like Luca (C5 in Table 2) and Xivo (similar to C4 in Table 2), are well-designed for this task. Objectively, they perform well in terms of design, interaction, and storytelling. The issue is, preschoolers have short attention spans, without any active learning consciousness. It is their basic characteristic. … I’ve heard that some children start to slack off once they’ve mastered the tool. They can quickly complete the reading task through repeatedly clicking the progress bar.” [P8]
@@ 4.3.2. Preschoolers Need the Sense of Authenticity and Constructiveness
The virtual interactive modality, without the sense of authenticity and constructiveness, was also reported by our participants as one major reason that current AI-based storytelling technologies couldn’t sustainably attract children’s attention and achieve the expected storytelling effect. They expressed that preschoolers were still in the stage of development of motor thinking, and their abstract thinking ability was not well developed. This thinking development stage determined that they needed more physical and real interactive processes with a sense of constructiveness, instead of the virtual way, as most existing intelligent tools used. They expressed that the virtual interactive way brought certain cognitive difficulties for preschoolers to form cognition in their minds. P15 used his son’s experience of playing with building blocks to explain this sense of authenticity and constructiveness:
@@
I purchased two expensive robots. In the beginning, there was a sense of novelty. My child interacted with them actively, saying things like ‘Egg, come with me; Egg, sing me a song; Egg, tell me a story.’ But he only played with them for three days. Currently, they are just expensive toys, sitting in the corner and serving no purpose. Children can’t establish an emotional connection with them. I think things that can be touched and felt will provide more satisfying experiences for children. [P15]
Additionally, our participants also considered systematic and contextual information as significant in constructing children’s cognition of the story, but lacked in AI-based storytelling. Consequently, it became challenging to use AI-based storytelling tools to support attractive storytelling for children. They stated that in traditional paper books-based storytelling, the storybooks often possessed coherent story structures and contextual situations. Throughout the reading process, children could perceive this contextual information and construct the stories in their minds through flipping the pages back and forth. In contrast, AI agents only presented the story plot at a specific moment, devoid of any contextual information. This limited their ability to develop children’s comprehensive understanding of the story, subsequently impacting their engagement and interests.
@@ 4.3.3. Preschoolers Need Interaction, Instead of a Correct Answer
Interaction and stimulating children’s active thinking were reported by our participants as a primary expectation of storytelling, even more important than learning knowledge. Through storytelling, they desired children would be guided to actively participate in the process of imagination, emotional expression, and critical thinking, which were helpful to enhance their creativity, empathy, and problem-solving skills. P13 showed us a vivid storytelling process he conducted with his 5-year-old daughter, which illustrated this interacting and inspiring purpose very well.
One day, I told my daughter the story of ‘monkey move corn’, in which the monkey initially held the corn. After seeing a peach, the monkey held the peach and threw the corn away. After seeing a watermelon, it then held the watermelon and threw the peach away. … In the end, the monkey didn’t catch anything. This classic story has many inherent truths, such as the importance of having a clear goal and being focused on it. However, I didn’t tell my daughter this moral directly. Instead, I asked her, ‘if you were the monkey, what would you do?’ To my surprise, my daughter responded with great insight. She said, ‘if I were the monkey, I might have three ways to solve the problem. First, I would bring a bigger bag to take everything. Second, I would pick corn today, peaches tomorrow, and watermelons the day after tomorrow. Third, I would call my friends to come with me, then we could gather corn, peaches, and watermelons together.’ [P13] ”
Based on this story, P13 and many other participants emphasized that meaningful interactive storytelling should aim to stimulate children’s active thinking, rather than simply imparting a fixed moral or lesson. However, most current AI-based storytelling tools pay primary attention to supporting correct question-answering, i.e., providing a ‘correct answer’ to children’s questions. They were completely unable to support their desired interactive storytelling. P9, for instance, described her perception of using Tmall Genie (a general conversational agent, similar to C6 in Table 2) to support storytelling practices for preschoolers:
When the child asks why the sky is blue, it gives an answer and then ends the conversation. This is of limited meaning to the child. It cannot inspire children’s interests and thinking. They may not understand it, and then their interest stops there. What preschoolers need is actually an interactive process, not just a correct answer.” [P9]
For this reason, some of our participants strongly rejected the tools with explicit educational purposes, because they might limit children’s thinking and self-reflection.
@@
Taken together, our participants illustrated a series of interactive challenges between the interactive ways existing AI-based storytelling tools are used and the interactions aligning with preschoolers’s cognitive and behavioral characteristics. They suggested that AI-based storytelling technologies should carefully consider preschoolers’s inherent characteristics and cognitive development level. Meanwhile, technologies should guide children’s active thinking, instead of telling them the correct answer or specific knowledge directly.
@@
The second category of challenges we identified was algorithmic challenges, mainly referring to challenges and dilemmas caused by AI algorithms embedded in storytelling tools. These challenges primarily arose from the limited AI algorithms in supporting child-centered, heuristic, personalized, and adaptive storytelling practices. As a result, the usability and effectiveness of these tools in facilitating practical, parent-expected intelligent storytelling for preschoolers were constrained.
@@ 4.4.1. Challenges in Supporting ‘Intelligent’ Interaction
Silly and limited-intelligent interaction was one major algorithmic challenge our participants experienced when they used AI-based tools, no matter children-tailored tools (e.g. C3, C4, C5 in Table 2) or general intelligent tools (e.g C6 in Table 2), to support the storytelling with their children. For children-tailored tools (e.g C1-C5 in Table 2), for instance, our participants commonly considered their intelligence was limited, because their interaction ways and content were notably inflexible, adhering to a fixed structure and only permitting content access and updates within the limitations set by the merchant. The experience of P1 using Little Genius (similar to C4 in Table 2) illustrated this limited intelligence very well: “its content is almost static, and the interaction remains unchanged. Whether you open it this year or next year, it remains the same. It lacks adaptability and fails to adjust according to the child’s developmental needs.
For general intelligent conversational agents based on massive corpora, such as C6 in Table 2, our participants considered that, although they performed commands well, their abilities in understanding children’s attention and supporting children-oriented dialogue and interaction were still very limited. P6’s experiences of using XiaoDu (C6 in Table 2) illustrated this limited intelligence very well:
@@
In particular, our participants highlighted that in real-life scenarios, children’s interests and questions were frequently diverse. However, current AI-based storytelling technologies are more like search engines. They could effectively answer questions that could be queried through search engines, but failed to answer those that couldn’t be queried. For example, P8 evaluated the use of Alpha Egg (C5 in Table 2):
The function is not particularly smart. It’s just a search engine. When you ask who is Osa, it searches online and reads the answer to you. Once your question can’t be searched online, then it fails to answer.” [P8]
Even with more intelligent generative AI technologies, parents in our study still expressed their concerns about their capacity to genuinely grasp and decipher preschoolers’s intentions and furnish responses. They expressed uncertainty about whether and how generative AI technologies could answer their children’s endless and personalized ”why” questions well, because many of these questions actually didn’t have standard answers.
Child’s questions are often very strange. They may ask why witches are female and not male, why witches have pointy noses, and why witches are thin instead of fat. Or ask why Harry Potter is named Harry Potter, why he rides a broomstick instead of a vacuum cleaner. There are no standard answers to these questions. How can machines answer these questions? If the machine asks the child, ‘do you think it’s right or wrong for the Yellow Robe Monster to steal someone’s wife and treat her especially well?’ and the child answers with their reasoning, how should the machine respond?” [P6]
@@ 4.4.2. Challenges in Supporting Children-Appropriated Interaction and Content Generation
Whether the AI tools could support children-appropriated interaction or generate children-appropriated content was another significant algorithm challenge reported by our participants. Our participants indicated that when they used tools such as C5 and C6 in Table 2 to answer children’s story-related questions, the generated answers were often serious and professional, not specifically tailored to children’s cognitive level and psychological age, which caused it difficult for Children to understand. P15’s experience of using XiaoAi to answer children’s questions illustrated this very well:
@@
A similar concern was also reported by participants who had used ChatGPT to answer children’s questions. They expressed that the answer generated by ChatGPT was not very suitable for preschoolers, particularly in terms of knowledge level, language structure, and presentation. As P8 said, “The content is often too long and professional. Many terms are unfamiliar to preschoolers. If it is used to talk to children, it may need to be shorter and simpler. I think this is a technological challenge. There are thousands of questions, and how can it answer them in the concise and appropriate language that children can understand?
@@ 4.4.3. Challenges in Supporting Personalized Interaction
Personalized interaction was reported by our participants as the basic interaction requirement of AI-based storytelling technologies. They emphasized that, in practice, each child’s characteristics, interests and questions were very different, “Some children are receptive, while others are more proactive” [P10], and “some children prefer to read by themselves, while others like you (the parents) to tell them” [P5]. These personalized features caused that, even if children saw the same story, their points of interest might be completely different. As P15 said, “with the same story, different children may focus on different things. Boys and girls think differently, and my child and her tablemate in kindergarten also think differently”.
In addition to differences among children, our participants emphasized that one child’s character, interests and focuses also dynamically change. Many participants told us that their children read the same story repeatedly, but their understanding and points of interest about the story content changed continuously as they grew older. However, although many existing AI-based storytelling technologies claimed that they supported personalized interaction and learning, in our participants’ view, current personalized interaction was very coarse-grained (e.g., “first graders may read these books, second graders may read these books” [P12]), which couldn’t fully support their expected personalized interaction with preschoolers. They emphasized that true personalization for preschoolers needed to consider various factors such as children’s real-time interests, needs, characteristics, personalities, etc., not just the grade and age. Meanwhile, these technologies didn’t provide the functions of automatically perceiving personalized characteristics of children. Instead, they often relied on parents manually inputting some basic characteristics of the children, such as age, gender, interest, etc., and based on these characteristics to conduct “coarse-grained personalization”. This made it more difficult for the so-called personalized features to become a reality in practical storytelling scenarios.
Some technologies employed recommendation algorithms to support personalized story and content recommendations. Yet, our participants considered that the current recommendation algorithms were almost incapable of providing human-centered personalized recommendations. They worked more like “the child listens to a story today, then the system will recommend the similar stories for the next few days” [P4]. P4 considered these recommendation algorithms as “greatly limiting the personalized development of preschoolers”. Our participants agreed that perceiving children’s personalized characteristics and interacting with children based on their personalized characteristics were highly challenging, even for very intelligent LLM-based tools such as ChatGPT. “Large models might know the overall patterns, but didn’t know the specific feature and need of each child”, as P7 said.

### v2

@@ 4.1.2. Parents’ Expectations to AI-based Storytelling Tools: “Being Parents’ Secret Weapon
4.2. Perceived Value of Using AI-Based Interactive Storytelling Technologies
When reflecting on the experience of using AI-based interactive storytelling tools to associate daily parent-child storytelling, our participants talked about some benefits of these technologies they perceived. For example, while all our participants agreed that interactive storytelling was a very important and beneficial parent-child activity for preschoolers, most of them expressed this conversation-rich interactive storytelling was full of practical challenges. As P9, a mother with a 6-year-old son, said, “I really want to interact with my son, but I don’t know what story content should be interacted. I even can’t understand the questions he asks me.” Some participants also noted that engaging in daily, high-quality interactive storytelling, as they expected, represented significant challenges in terms of time, energy, and physical stamina. This was particularly true for today’s young parents who were often immersed in fast-paced and demanding work.
@@ 4.2. Perceived Value of Using AI-Based Interactive Storytelling Technologies
At the same time, many of our participants stated that compared to plush toys and the like, these AI-based tools were relatively “better” toys for children, because they could bring “novelties and a sense of technology” [P8, P13] to children, which parents couldn’t bring. Most existing storytelling tools had some children-tailored interactive designs, such as immersive storytelling blending multiple modalities (“Children could touch, hear, see, and talk with it.” [P6]), and vivid presentations (“The sound of the storyteller is very pleasant and full of emotions and varied intonations, and the animals can move and make sounds” [P1]), which worked well to attract children to participate in the interaction, improving their storytelling experience significantly.
In addition, most of these technologies integrated more or less learning resources, or had some child-tailored interactive designs, which allowed children to learn something or develop certain abilities while playing. As P8, a mother of 6-year-old son, said: “You always need to buy something for the child, right? Do you buy a bunch of Ultraman or cloth dolls, or these intelligent things? No matter how the practical effect is, it provides the possible opportunities to cultivate the creative thinking abilities of children, or learn something when playing.” [P8]
Despite these perceived benefits, our participants considered that most of the existing storytelling tools they had bought and used were hard to consistently engage preschoolers and often struggled to achieve the intelligent features claimed by the products. The practical effectiveness and efficiency in supporting quality storytelling were limited. Our participants reported several challenges with these tools that hindered their practical effectiveness. We generally categorized these challenges into two types - interactive and algorithmic, detailed in the following two sections.
4.3. Interactive Challenges in Supporting Storytelling Aligned with Preschoolers’ Cognitive Characteristics
Interactive challenges mainly refer to challenges and dilemmas caused by interactive ways and interfaces existing AI-based storytelling tools used. Our participants stated that existing AI-based storytelling tools they had used didn’t consider preschool-age children’s cognitive characteristics and their abilities to accept and use these technologies, which hindered the tools’ practical effectiveness.
@@ 4.3.1. Preschoolers are Difficult to Use AI Tools Proactively, Independently, and Consistently
According to our participants, preschoolers often exhibited short attention spans, lacked self-learning and self-management abilities, etc. These characteristics made it hard for them to use AI-based tools actively, independently, and consistently. Most of our participants thought it was unrealistic to expect preschoolers to use these tools to read a storybook independently. Instead, most tools require parent-engaged usage processes. For example, the use of an interactive talking pen (C1 in Table 2) required parents to “take out the pen, turn it on, find the corresponding book, and flip to the specific story [P8]” and then the children could use it to read a story. “Without parents’ operations and guidance, preschoolers are unlikely to proactively use these tools to read the story” [P8]. P15, who had purchased two conversational agents tailored for kids (Similar to C5 in Table 2) for his 3-year-old daughter and 10-year-old son, also expressed that ”today’s children have too many interesting things to do. Without parents’ engagement, they may only use them (conversational agents) for a short time out of curiosity and then turn to do other things [P15].”
Further, preschoolers’ short attention spans made it challenging to consistently and sustainably use any tools with focused attention. Whether and how parents were involved affected the practical effectiveness of tools. Many parents in our study expressed that there was a notable difference in effectiveness between children using these tools independently and when accompanied by parents. P5, a mother of a 6-year-old daughter, considered that, “if parents are not involved, even for the best product, it is difficult to sustain children’s interest and attention, and motivate them to read a story with it for 15-20 minutes steadily”. P8, who had purchased many AI-based storytelling tools for her son, shared her experience with these tools.
Our (parents’) involvement or not has a completely different using effect. I can let these tools tell stories. Tools, like Luca (C5 in Table 2) and Xivo (similar to C4 in Table 2), are well-designed for this task. Objectively, they perform well in terms of design, interaction, and storytelling. The issue is, preschoolers have short attention spans, without any active learning consciousness. It is their basic characteristic. … I’ve heard that some children start to slack off once they’ve mastered the tool. They can quickly complete the reading task by repeatedly clicking the progress bar.” [P8]
@@ 4.3.2. Preschoolers Need the Sense of Authenticity and Constructiveness
The virtual interactive modality, without the sense of authenticity and constructiveness, was also reported by our participants as one major reason that current AI-based storytelling technologies couldn’t sustainably attract children’s attention and achieve the expected storytelling effect. They expressed that preschoolers were still in the stage of development of motor thinking, and their abstract thinking ability was not well developed. This thinking development stage determined that they needed more physical and real interactive processes with a sense of constructiveness, instead of the virtual way, as most existing intelligent tools used. P15 used his son’s experience of playing with building blocks to explain this sense of authenticity and constructiveness:
@@
In the beginning, there was a sense of novelty. My child interacted with them (C5 in Table 2) actively, saying things like ‘Egg, come with me; Egg, sing me a song; Egg, tell me a story.’ But he only played with them for three days. Currently, they are just expensive toys, sitting in the corner. Children can’t establish an emotional connection with them. I think things that can be touched and felt will provide more satisfying experiences for children. [P15]
Additionally, our participants also considered systematic and contextual information as significant in constructing children’s cognition of the stories, but lacked in AI-based storytelling scenarios. They stated that in traditional paper books-based storytelling, the storybooks often possessed coherent story structures and contextual situations. Throughout the reading process, children could perceive this contextual information and construct the stories in their minds through flipping the pages back and forth. However, AI agents only presented the story plot at a specific moment, devoid of any contextual information. This limited their ability to develop children’s comprehensive understanding of the story, subsequently impacting their engagement and interests.
@@ 4.3.3. Preschoolers Need Interaction, Instead of a Correct Answer
Interaction and stimulating children’s active thinking were reported by our participants as one primary expectation of interactive storytelling, even more important than learning knowledge. Through interactive storytelling, they desired children could actively participate in the process of imagination, emotional expression, and critical thinking, which were helpful to enhance their creativity, empathy, and problem-solving skills. P13, for instance, showed us an interactive storytelling process he conducted with his 5-year-old daughter, which illustrated this interacting process.
One day, I told my daughter the story ‘Monkey move corn’, in which the monkey initially held the corn. After seeing a peach, the monkey held the peach and threw the corn away. After seeing a watermelon, it then held the watermelon and threw the peach away. … After seeing a rabbit, it chased the rabbit and threw the watermelon away again. In the end, the monkey didn’t catch anything. This classic story has many inherent truths, such as the importance of having a clear goal and being focused on it. However, I didn’t tell my daughter this moral directly. Instead, I asked her, ‘if you were the monkey, what would you do?’ To my surprise, my daughter responded with great insight. She said, ‘if I were the monkey, I might have three ways to solve the problem. First, I would bring a bigger bag to take everything. Second, I would pick corn today, peaches tomorrow, and watermelons the day after tomorrow. Third, I would call my friends to come with me, then we could gather corn, peaches, and watermelons together.’ [P13] ”
Based on this story, P13 and many other participants emphasized that meaningful interactions should aim to stimulate children’s active thinking, rather than simply imparting a fixed moral or lesson. However, most current AI-based storytelling tools pay primary attention to supporting correct question-answering, i.e., providing a ‘correct answer’ to children’s questions. They were completely unable to support their desired interactive storytelling. For this reason, some of our participants strongly rejected the tools with explicit educational purposes, because they might limit children’s thinking and self-reflection.
@@
Given these interactive challenges, our participants suggested that AI-based interactive storytelling technologies should carefully consider preschoolers’s inherent characteristics and cognitive development level. Meanwhile, technologies should guide children’s active thinking, instead of just telling them the correct answer or specific knowledge.
@@
Algorithmic challenges primarily arose from the limited ability of AI algorithms to support child-centered, heuristic, personalized, and adaptive storytelling practices, which further caused the limited usability and effectiveness of existing AI-empowered tools in facilitating practical, parent-expected intelligent storytelling for preschoolers.
@@ 4.4.1. Challenges in Supporting ‘Intelligent’ Interaction
Silly and limited-intelligent interaction was one major algorithmic challenge our participants experienced when they used AI-based tools, no matter children-tailored tools (e.g. C3, C4, C5 in Table 2) or general intelligent tools (e.g C6 in Table 2). For children-tailored tools, our participants commonly considered their intelligence was limited, and interaction ways and content were notably inflexible, adhering to a fixed structure and only permitting content access and updates within the limitations set by the merchant. The experience of P1 using Little Genius (similar to C4 in Table 2) illustrated this limited intelligence very well: “its content is almost static, and the interaction remains unchanged. Whether you open it this year or next year, it remains the same. It lacks adaptability and fails to adjust according to the child’s developmental needs.
For general intelligent conversational agents based on massive corpora, such as C6 in Table 2, our participants considered that, although they performed commands well, their abilities in supporting children-oriented dialogue and interaction were still very limited. P6’s experiences of using XiaoDu (C6 in Table 2) illustrated this limited intelligence very well:
@@
In particular, our participants highlighted that in real-life scenarios, children’s interests and questions were frequently diverse. Yet, current AI-based storytelling technologies were more like search engines. They could effectively answer questions that could be queried through search engines, but failed to answer those that couldn’t be queried. As P8 evaluated the function of Alpha Egg (C5 in Table 2): “It’s just a search engine. When you ask who is Osa, it searches online and reads the answer to you. Once your question can’t be searched online, then it fails to answer.” [P8] Even with more intelligent generative AI technologies, parents in our study still expressed their concerns about their capacity to genuinely grasp and decipher preschoolers’s intentions and furnish responses. They expressed concerns about whether and how generative AI technologies could answer children’s endless and personalized “why” questions well, because many of these questions actually didn’t have standard answers. As P6 said,
Child’s questions are often very strange. They may ask why witches are female, why witches have pointy noses, why witches are thin instead of fat, why Harry Potter is named Harry Potter, why he rides a broomstick instead of a vacuum cleaner… There are no standard answers to these questions. How can machines answer these questions? ” [P6]
@@ 4.4.2. Challenges in Supporting Children-Appropriated Interaction and Content Generation
Whether AI tools could support children-appropriated interaction or generate children-appropriated content was another significant algorithm challenge reported by our participants. Our participants indicated that when they used tools such as C5 and C6 in Table 2 to answer children’s story-related questions, the generated answers were often serious and professional, not specifically tailored to children’s cognitive level and psychological age, which caused it difficult for Children to understand. P15’s experience of using XiaoAi to answer children’s questions illustrated this very well:
@@
A similar concern was also reported by participants who had used ChatGPT to answer children’s questions. They expressed that the answer generated by ChatGPT was not very suitable for preschoolers, particularly in terms of knowledge level, language structure, and presentation. As P8 said, “The content is often too long and professional. Many terms are unfamiliar to preschoolers. If it was used to interact with children, it needed to be shorter and simpler. I think this is a big technological challenge. There are thousands of questions, and how can it answer them in the concise and appropriate language that children can understand?
@@ 4.4.3. Challenges in Supporting Personalized Interaction
Personalized interaction was reported by our participants as the basic interaction requirement of AI-based storytelling technologies. They emphasized that, in practice, each child’s characteristics, interests and questions were different, “some children are receptive, while others are more proactive” [P10], and “some children prefer to read by themselves, while others like you (the parents) to read for and interact with them” [P5]. These personalized features caused, even if children saw the same story, their points of interest might be completely different. “With the same story, different children may focus on different things. Boys and girls think differently, my child and her tablemate in kindergarten also think differently” [P15]. Further, our participants emphasized that one child’s character, interests and focuses also dynamically change. Many participants told us that their children read the same story repeatedly, but their understanding and points of interest about the story content changed continuously as they grew older.
However, although many existing AI-based storytelling technologies claimed that they supported personalized interaction and learning, in our participants’ view, current personalized interaction was very coarse-grained (e.g., “first graders may read these books, second graders may read these books” [P12]), and couldn’t fully support their expected personalized interaction. They emphasized that true personalization for preschoolers needed to consider various factors such as children’s real-time interests, needs, characteristics, personalities, etc., not just the grade and age. Meanwhile, these technologies didn’t provide the functions of automatically perceiving personalized characteristics of children. Instead, they often relied on parents manually inputting some basic characteristics of the children and based on these characteristics to conduct “coarse-grained personalization”. This made it more difficult for the personalized features to become a reality in practical interactive storytelling scenarios.
Even some technologies employed recommendation algorithms to support personalized story and content recommendations, our participants considered them as “almost incapable of providing human-centered personalized recommendations” [P8]. They thought these tools worked more like “the child reads a story today, then the system recommends the similar stories for the next few days” [P4], which might “greatly limit the personalized development of preschoolers” [P4].